### v1

# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

Abstract
In the field of Prognostics and Health Management (PHM), recent years have witnessed a significant surge in the application of machine learning (ML). Despite this growth, the field grapples with a lack of unified guidelines and systematic approaches for effectively implementing these ML techniques and comprehensive analysis regarding industrial open-source data across varied scenarios. To address these gaps, this paper provides a comprehensive review of machine learning approaches for diagnostics and prognostics of industrial systems using open-source datasets from PHM Data Challenge Competitions held between 2018 and 2023 by PHM Society and IEEE Reliability Society and summarizes a unified ML framework. This review systematically categorizes and scrutinizes the problems, challenges, methodologies, and advancements demonstrated in these competitions, highlighting the evolving role of both conventional machine learning and deep learning in tackling complex industrial tasks related to detection, diagnosis, assessment, and prognosis. Moreover, this paper delves into the common challenges in PHM data challenge competitions by emphasizing both data-related and model-related issues and summarizes the solutions that have been employed to address these challenges. Finally, we identify key themes and potential directions for future research, providing opportunities and prospects for ML further development in PHM.
1 Introduction
In the contemporary landscape of Industry 4.0, the emphasis on the reliability, efficiency, and longevity of industrial systems has become crucial [34]. Against this backdrop, Prognostics and Health Management (PHM) has crystallized as an indispensable discipline, that integrates the detection, diagnosis, assessment, and prognosis of system failures addressing the growing need for proactive system health management [82]. This integrative approach can bolster the reliability and safety of industrial engineering systems, with minimal unplanned downtimes and reduced maintenance costs.
The rise of the Internet of Things [62], big data analytics [74], cyber-physical systems [35], machine learning (ML) [21], deep learning (DL) [57], and industrial artificial intelligence [54, 36] has paved the way for a transformative shift in PHM. Historically, PHM strategies largely relied on physical-based methods, which derived their strength from profound insights into system physics, material properties, and failure mechanisms. However, these strategies often struggled with scalability, adaptability, and handling the vast variability and uncertainties inherent in real-world operations. Data-driven approaches, in contrast, underpinned by machine learning algorithms, offer the ability to model complex systems, identify intricate patterns, and diagnose and predict failures with high accuracy. This evolution underscores the potential of data-driven models to deliver precise and adaptable diagnostic and prognostic solutions, even in the absence of explicit physical models.
Over the last decade, organizations such as the PHM Society and the IEEE Reliability Society have played pivotal roles in fostering innovation, research, and collaboration in the PHM field. One notable initiative is the PHM data challenge competition, which stands as a testament to the growing emphasis on real-world, data-centric PHM solutions. By providing participants with genuine industrial datasets and posing real-world challenges, this competition seeks to accelerate the development and validation of cutting-edge PHM methodologies, bridging the gap between academia and industry. In light of the recent surge in data-driven methodologies, particularly machine learning, and deep learning, in PHM, it becomes essential to revisit and scrutinize research algorithms and methodologies derived from the PHM data challenge competitions and their datasets. Systematically reviewing the challenges presented by industrial data across various sectors with methodologies and analytics can foster and enhance a deeper understanding of data-driven approaches. Such endeavors not only propel the evolution of data-centric techniques, machine learning, and artificial intelligence (AI) within the PHM domain but also accentuate advancements in the broader landscape of industrial data analytics and industrial artificial intelligence.
This study aims to conduct a comprehensive problem-challenge-solution-oriented review within the domain of PHM Data Challenge. Specifically, this study summarizes the problems and challenges presented in nine PHM data challenge competitions, elucidating the algorithms, methods, and analytical techniques employed to address them. Further, this study encapsulates the common challenges associated with industrial data, underscoring specific issues related to data issues, and modeling in data-driven approaches. In pursuit of the objective, approximately sixty research papers were reviewed, encompassing both competition-winning contributions and subsequent exploratory research undertaken post-competition.
The rest of this paper is organized as follows: Section 2 provides an overview of PHM data challenge competitions and underscores the principal research tasks within PHM. Section 3 and Section 4 respectively introduce the prevalent challenges associated with two pivotal tasks: detection & diagnosis, and assessment & prognosis, subsequently detailing the solutions presented in various PHM data competitions individually. Section 5 critically summarizes and examines common challenges in the PHM domain through the lens of data-driven methodologies, addressing concerns related to data issues, and model issues. Section 6 provides a forward-looking perspective on the enhanced integration of data-driven techniques, machine learning, and AI in future PHM endeavors. Section 7 concludes the paper, highlighting its findings and contributions.
2 Background
In this section, we begin with an overview of the PHM data challenge competition. Next, we outline the core research tasks in the PHM field and provide their respective explanations. We then summarize the issues and predominant challenges posed by various competitions and review the data-driven approaches used in the competitions.
2.1 Overview of PHM Data Challenge Competitions
From 2018 to 2023, PHM Society and IEEE Reliability Society have initiated nine PHM data challenge competitions which reflect challenges associated with analyzing industrial data across diverse industrial sectors. The main purpose behind these competitions is to amplify focus and collaborative efforts from both academic and industrial realms to address real-world challenges. The topics of these competitions are commendably extensive, encompassing diverse industrial entities such as Ion Mill Etching Tools, Filtration Systems, Manufacturing Production Lines, Turbofan Engines, Printed Circuit Boards, Rock Drills, Spacecraft Propulsion Systems, and Gearbox, among others. Furthermore, the range of problems posed encapsulates the main tasks in the PHM domain including detection, diagnosis, assessment, and prognosis. These tasks align with PHM’s ultimate goal: to accurately forecast system health, degradation, and eventual failure, thereby improving system reliability, safety, and operational efficiency. For ease of discussing different competitions, we’ve adopted a condensed naming convention for the PHM data challenge competition: ”YEAR ORGANIZATION” indicates the respective year and organizing body of the challenge. Among the challenges discussed, one is organized by IEEE Reliability Society, while the remaining eight are conducted by PHM Society. It’s noteworthy that the PHM Society conducts an annual conference in North America, an Asia-Pacific conference in odd years, and a European conference in even years. We use the abbreviations ”PHM NA”, ”PHM AP”, and ”PHM EU” to represent the competitions held in North America, Asia-Pacific, and Europe, respectively. For instance, ”2018 PHM NA” refers to the challenge held in North America by the PHM Society in 2018 while ”2023 IEEE” represents the competition hosted by the IEEE Reliability Society in 2023. Table 1 presents the systems, research tasks, and common challenges of the nine data competitions. For an in-depth overview of the PHM data challenge competitions and their associated datasets, please refer to the appendix.
Competition Industrial Systems Research Task
2018 PHM NA Ion Mill Etching System Detection & Diagnosis & Assessment & Prognosis
2019 PHM NA Fatigue Crack Assessment & Prognosis
2020 PHM EU Filtration System Prognosis
2021 PHM EU Manufacturing Production Line Detection & Diagnosis
2021 PHM NA Turbofan Engine Prognosis
2022 PHM EU Printed Circuit Board Detection & Diagnosis
2022 PHM NA Rock Drill Detection & Diagnosis
2023 PHM AP Spacecraft Propulsion System Detection & Diagnosis
2023 IEEE Gearbox Detection & Diagnosis
Common Challenges: Data Imbalance, Domain Shift, Multiple Fault Modes, Unknown Fault, Limited Data Availability
Table 1: Overview of PHM Data Challenge Competitions from 2018 to 2023
2.2 Core Research Tasks within PHM
In the realm of PHM, there are four pivotal research tasks: detection, diagnosis, assessment, and prognosis.
Detection: The foundational step of PHM. It entails continuous monitoring of equipment and systems to identify any anomalies or deviations from expected performance. Notably, the primary aim of this phase is anomaly identification (detection) rather than root cause analysis.
Diagnosis: Upon detecting anomalies, the diagnostic phase delves deeper to determine failure types or failure modes and may find out the root causes of the problem, enabling targeted corrective actions in a timely manner. In diagnostic scenarios, detected failures often need to be classified into specific failure types.
Assessment: Following the identification of the root cause, the assessment phase evaluates the current operational status and performance of the system. Leveraging either historical data or recent machinery behavior, this phase evaluates potential risks or assesses the health status of the system in its present condition.
Prognosis: Representing the most forward-looking phase, prognosis leverages both current and historical data to forecast the future health of a system or the residual life of a machine. Commonly, this is referred to as predicting the Remaining Useful Life (RUL). This pivotal phase provides estimates on potential system or machinery failure timelines, thereby facilitating proactive maintenance planning. To achieve this, data-driven approaches, including machine learning and artificial intelligence, are often employed to construct predictive models for future prognosis.
Figure 1: Research Tasks in PHM Data Challenge Competitions.
Reflecting on the nine data competitions spanning the last six years, Figure 1 illustrates the associated research tasks. Our analysis yielded the following insights:
Integration of Fault Detection and Diagnosis: Fault detection and diagnosis frequently co-occur in research tasks. While fault detection primarily identifies discrepancies or anomalies compared with expected performance, it doesn’t delve into their root causes or subsequent maintenance implications. Following fault detection, diagnosis intensively examines the root causes of the faults, categorizing them by distinct fault types. This categorization aids in guiding subsequent maintenance operations.
Synergy between Health Assessment and Prognosis: There exists an intrinsic link between tasks centered on health assessment and those focused on prognosis. While health assessment provides the current state of system health, prognostic tasks often predict future states based on the current assessment. Notably, even in competitions or tasks primarily targeting prognosis, emphasis is occasionally placed initially on health assessment methodologies. This precedence stems from the understanding that a robust health assessment forms the foundation for efficacious prognostic strategies. Recognizing a system’s current health state equips researchers with a more potent toolkit to tackle prognostic challenges.
2.3 Comprehensive Review of Data-Driven Approaches in Recent PHM Data Challenge Competitions
Figure 2: A Unified Machine Learning Framework in Prognostics and Health Management Domain.
After conducting an in-depth review of data competitions from the last six years, we have summarized a unified machine learning framework that concludes machine learning approaches for PHM data competitions during this period. As shown in Figure 2, it encompasses five primary components: Data Collection, Data Processing, Data Visualization, Conventional Machine Learning & Deep Learning, and Model Interpretability.
Data Collection: Data collection is the foundation of these open-source data challenge competitions. Companies across various industries contribute datasets with the objective of harnessing data-driven strategies to address intricate challenges in PHM. Typically, these industrial datasets encompass a diverse range of data types, including but not limited to time-series data, tabular data, images, sensor readings, and simulation data.
Data Processing: A pivotal stage before ML model development is data processing, crucial for enhancing data quality and ensuring effective model training. This stage can be further subdivided into two aspects: data preprocessing and feature engineering. Data preprocessing addresses raw data challenges, typically including missing data, noise, outliers, data imbalance, and scaling issues. Techniques like imputation, denoising, outlier detection, resampling, and normalization/standardization are commonly applied. Feature engineering follows, refining data representation post-preprocessing. This step often uncovers hidden insights and deepens understanding of data, therefore significantly enhancing model performance and predictive capabilities in the later stage.
Data Visualization: Data visualization is another important aspect of the machine learning process. It involves transforming data into intuitive graphical representations, such as graphs or charts. Researchers can use these charts to better observe trends, patterns, or outliers in data, which further help people have a better understanding of data and generate some useful insights. In the PHM domain, effective visualization can help initial data exploration and analysis, and accelerate data processing and ML modeling development.
Conventional Machine Learning & Deep Learning Techniques: Post data processing, various conventional machine learning, and deep learning methods are developed to solve the competition problems using useful and usable data, as shown in Figure 2. This segment includes ”model training and testing” and ”model prediction and classification”. During the training and testing phase. Observed techniques in these competitions are diverse, including Tree-Based Methods, CNNs, RNNs and their variants, SVMs, Unsupervised Learning, Ensemble Learning, Similarity-Based Methods, Rule-Based Methods, Transfer Learning, Domain Adaptation, etc. Once the machine learning models are trained and tested well, they become invaluable tools for various PHM applications which include detecting abnormal performance or faults, classifying different root cause faults, assessing the current health state of systems, or predicting the remaining useful life of components. These applications are not only critical in meeting the diverse requirements of PHM data challenge competitions but also demonstrate the adaptability and effectiveness of the various machine learning and deep learning approaches utilized.
Model Interpretability: Except for the pursuit of accuracy in predictions and classifications, model interpretability is becoming more and more important. Model interpretability refers to interpreting machine learning model outputs [52] which helps humans understand the ’why’ behind a model’s predictions, facilitating collaboration and more informed decision-making. Many explainable artificial intelligence (XAI) methods such as Local Interpretable Model-Agnostic Explanations (LIME) [58] and SHapley Additive exPlanations (SHAP) [48] are utilized in the PHM data challenge competitions to explain model outputs in PHM [66].
In Section 3 and Section 4, our analysis will categorize and delve deeper into the dual themes of ”Detection & Diagnosis” and ”Assessment & Prognosis”.
3 Detection & Diagnosis
Fault detection and diagnosis are essential components of PHM, tasked with the identification of anomalies and determining failure modes in a timely manner, thereby playing a pivotal role in enhancing system safety and reliability, guaranteeing prompt maintenance interventions, and reducing downtime costs. In this section, we provide a comprehensive overview of the detection and diagnosis challenges encountered in the competitions, as illustrated in Table 2. We sequentially introduce the fundamental competition information, including background, objectives, challenges, and datasets, highlighting the innovative approaches to fault detection and diagnosis problems. A summarization of the competition problems and their datasets is encapsulated in Table 3. Furthermore, we provided a comprehensive summary of the methodologies utilized for detection and diagnosis within the PHM data challenge competition at the end of this section, as outlined in Table 4.
3.1 Challenges
Table 2 provides a detailed summary of the detection and diagnosis challenges specific to each competition. While each competition faces unique challenges due to its distinct industry background and objectives, a common issue they all share is the need to accurately classify multiple fault modes. Apart from the common issue, notably, the PHM EU competition in 2021 and the PHM AP in 2023 are both constrained by their small datasets, with the former comprising only 70 training and 29 test samples, and the latter having 177 training and 46 test samples. This paucity of data presents substantial obstacles for training machine learning models due to limited information from data. Additionally, the class imbalance present in the datasets of the 2021 and 2022 PHM EU competitions shows a significant imbalance with a higher number of healthy experiments compared to the faulty ones, thus impeding the models’ ability to detect rare fault conditions and challenging their robustness and generalization potential. In the meantime, these datasets also faced the obstacle of missing data, necessitating imputation techniques for robust data handling. The 2022 PHM NA competition confronts a domain-shift problem, where varying operational conditions and data from different assets result in inconsistent signal distributions, which complicates the task of accurate fault classification. Lastly, the 2023 PHM AP competition includes unexpected fault modes in the testing data, absent from the training set, reflecting the unpredictability of real-world industrial settings and underscoring the necessity for constructing robust and generalized models.
Unique Challenges
2021 PHM EU
Missing Data & Noise
Data Imbalance
Limited Data
2022 PHM EU
Missing Data
Data Imbalance
Component-level Prediction
2022 PHM NA Domain Shift
2023 PHM AP
Unknown Fault
Limited Data
2023 IEEE NA
Common Issues: Multiple Fault Modes
Table 2: Overview of Detection and Diagnosis Challenges in PHM Data Competitions
2021 PHM EU 2022 PHM EU 2022 PHM NA 2023 PHM AP 2023 IEEE
System Manufacturing Production Line Printed Circuit Boards Rock Drill Spacecraft Propulsion System Gearbox
Failure Mode 8 1+1+2 10 3 4
Asset NA NA 8 4 NA
Operating Condition 2 1+1+1 3 1 2
Data Type Time Series Tabular Pressure Signals Pressure Signals Vibration Signals
Volume of Data Limited Large Large Limited Large
(70 Train, 29 Test) (SPI:5985382, AOI:31617) (37229 Train, 16396 Test) (177 Train, 46 Test) (Total 50000)
Sampling Rate 0.1 Hz NA 50 kHz 1 kHz 10 kHz
Sampling Interval 1-3 hours NA NA 1.2 s 5 minutes
Table 3: Overview of Detection and Diagnosis Problems in PHM Data Competitions
3.2 2021 PHM EU (Manufacturing Production Line)
In a collaborative effort with the Swiss Centre for Electronics and Microtechnology (CSEM), a data challenge competition was introduced, offering a dataset derived from a real-world industrial manufacturing line dedicated to testing electrical fuses. The objective of this competition is to perform fault identification and classification, root cause analysis, rapid prediction, and system operation parameter identification. The dataset showed eight unique system failure modes, replicated under two distinct operating conditions. However, its scale was rather restricted, comprising merely 70 training and 29 testing samples. The dominant challenge with the training data was its class imbalance—most samples depicted a healthy condition. This scenario, coupled with the limited volume of data, presented challenges to the data-driven approach.
Amidst this backdrop, the winning solution was a notable combination of decision tree algorithms and a propagation system [12]. While the decision trees zeroed in on observational diagnoses, the propagation system skillfully tackled chronological intricacies by infusing a Kalman-style filter. This fusion bolstered the reliability of the outcomes. To address the prominent data imbalance, the SMOTE (Synthetic Minority Oversampling Technique) method was utilized. In a related exploration, Ince et al. [24] incorporated the Leave One Feature Out Importance (LOFO-Importance) package for discerning essential features. Subsequently, they applied linear discriminant analysis for dimensionality reduction. During the modeling phase, different ML methods were employed — gradient boosting algorithms such as XGBoost and LightGBM, LDA classifier, and Gaussian process classifier. Especially for gradient boosting, they used Genetic Algorithms (GA) to optimize hyperparameters. A pivotal observation was XGBoost’s superior performance over the other algorithms. Diverging in their approach, Aimiyekagbon et al. [2] showcased a rule-based diagnostic technique, comparing its performance with that of Decision Trees and Random Forest methods. On the other hand, Aydemir et al. [4] proposed a regularized LSTM for sifting through vital features and then leveraged an ensemble of binary LSTM classifiers for fault detection and classification.
Moreover, the tree-based algorithm XGBoost found favor not just in the aforementioned studies but also in three other distinct research papers [3, 55, 71] outside the competition. The distinctions among these studies were as follows: For instance, Tian et al. [71] relied on a feature importance ranking (FIR) method, targeting enhanced performance and simplification in intricate industrial classification scenarios, whereas Ramezani et al. [55] zeroed in on the challenge of missing value imputation, harnessing Partial Least Squares (PLS-MV). A significant contribution of Ramezani’s framework was its explainability, achieved by pinpointing key performance indicators for each fault family using the SHAP method [48]. In conclusion, tree-based methods can significantly enhance fault diagnosis, thereby boosting the efficiency and safety of manufacturing production lines.
3.3 2022 PHM EU (Printed Circuit Boards)
The 2022 European PHME Data Challenge, hosted in collaboration with Bitron Spa, spotlights a classification issue within an actual industrial Printed Circuit Board (PCB) production line. The challenge’s objectives span three pivotal sub-tasks: predicting Automatic Optical Inspection (AOI) defect detection based on Solder Paste Inspection (SPI) data, predicting human-made visual inspection labels (OperatorLabel), and finally, predicting the human-assigned repair label (RepairLabel). It’s noteworthy that, among the nine datasets discussed in this paper, this dataset stands unique. While others are time series data, this one is parametric data. Due to the extensive dataset, tree-based and neural network-based machine learning methods usually come to the fore.
Delving into the competition’s outcomes, Gaffet’s team clinched 1st place with an XGBoost method based on encoding and feature engineering [18]. Notably, they further harnessed the SHAP method [48] for model interpretation. Echoing a similar strategy, Taco et al. [68] leveraged two tree-based gradient boosting algorithms, namely LightGBM and XGBoost. Their approach centered on solving classification problems by extracting significant statistical data during the feature engineering phase. The importance of feature engineering was further emphasized by Tang et al. [70]. They introduced a novel statistical feature extraction method coupled with a PinNumber-based technique. This method aimed to compress pin-level data into component-level information. When predicting automatic inspection defects, they integrated a neural network model, factoring in feeding imbalance control to navigate data imbalance challenges. Their toolkit also boasted a random forest model for both human inspection and repair predictions. Schmidt’s team [59], though paralleling Tang et al. in several aspects, differentiated their approach in ML model construction. Apart from their extensive focus on feature engineering, they applied a multi-layer perceptron neural network for defect predictions in automated inspections. For human inspection outcomes, a random forest algorithm was their choice, while decision trees were favored for predicting human repairs.
Outside of competition, Mirzaei’s team delved deep into the challenge of imbalanced data [51]. Their solution entailed a data-level technique that integrated Recursive Feature Elimination (RFE) for feature selection and oversampling methods. This blend ensured balanced representation for minority classes, enhancing the performance of multiple ML algorithms, from Decision Trees and Random Forests to SVMs and 1dCNNs. In a somewhat divergent approach, Lee et al. [37] unveiled a new quality management paradigm termed Stream-of-Quality (SoQ) tailored for multi-stage manufacturing processes. By leveraging this dataset, they showcased the potency of their methodology. The implications of their work are vast, offering promising avenues for refining industrial AI algorithms and methodologies in a systematic manner.
3.4 2022 PHM NA (Rock Drill)
Rock drills play a pivotal role in sectors like mining, tunneling, and construction. Due to the potential economic and human costs from work interruptions caused by rock drill faults, ensuring accurate fault diagnosis is imperative. This data challenge tackles the intricate task of fault classification within various rock drill configurations. Challenges arise due to the domain-shift phenomenon, stemming from data sourced from diverse rock drill machines, which results in heterogeneous signal distributions that potentially impair classification accuracy. The dataset [27] encompasses ten distinct fault modes and one health mode, with training, validation, and testing data sizes of 34,045, 3,184, and 16,396, respectively.
In the competition, the top accolade was clinched by Oh et al. [53], who championed a hybrid strategy, melding data-driven techniques with signal-processing methods. Their approach harnessed domain adaptation, metric learning, and pseudo-label-based deep learning to construct an ensemble deep learning model for comprehensive fault classification. For samples that posed challenges for deep learning, they deployed signal processing methods like Dynamic Time Warping (DTW), and Cross-correlation, and used SVM for supervised learning. The runner-up solution introduced a data-cropping technique, employing a convolutional neural network (CNN) as a pivotal feature extractor to bridge data length discrepancies. Kim et al. [31] innovatively addressed the domain-shift issue through a domain-adaptation-based scheme that harnessed a domain adversarial learning neural network for extracting domain invariant features while using maximum mean discrepancy (MMD) minimization for bridging distribution discrepancy, and a soft voting ensemble to reduce model uncertainty. Moreover, Minami and team proposed an ensemble approach, appending specialized models onto a baseline (or initial) model. This ensemble incorporated domain adaptation strategies to accommodate domain fluctuations. The baseline model employed conventional ML algorithms like SVM, Random Forest, and XGBoost for whole multi-class classification, whereas the specialized sub-models, leveraging CNN for feature extraction and classification, concentrated on binary classifications targeting specific classes that poorly performed on the baseline model [50].
Outside the competition, two notable data-driven strategies emerged. Ling’s team [43] deployed an end-to-end fault classification framework derived from X-Vectors [63], achieving integrated optimization of both feature extraction and classification phases. Taco et al. [69] proposed a novel deep forest algorithm that fused multi-grained scanning for feature extraction and a cascade forest for layered predictions to perform failure model classification. In summary, all aforementioned methods emphasize the importance of tackling the domain shift challenge. Utilizing domain adaptation and machine learning techniques can enhance the accuracy of fault detection and diagnosis in rock drills, ensuring safer and more reliable operations.
Methodology
Deep Learning Conventional Machine Learning Feature Engineering
2021 PHM EU
Regularized LSTM
Ensemble LSTM
XGBoost, Decision Trees
Random Forest
LDA, Rule-based Method
Gaussian Process
FIR
PLS-MV
FCM
SMOTE
2022 PHM EU 1D-CNN
Decision Trees, Random Forest
XGBoost, LightGBM
SVM, MLP
RFE
Statistical Feature Extraction
2022 PHM NA
Domain Adaption
DANN, X-Vectors
Metric Learning
Pseudo Label Technique
Ensemble Learning
XGBoost
SVM
Deep Forest Algorithm
MMD
RFE
DTW
2023 PHM AP NA
Similarity-based Method
K-means clustering, KNN
Decision Trees, XGBoost
Rule-based method
Ensemble Learning
Physical Feature Extraction
DTW
2023 IEEE
ROCKET, LSTM-FCN
1D-CNN with ResNet
Deep Residual Network
Residual-based CNN
Ensemble Learning
Data Augmentation
Data Regularization
STFT
Table 4: Overview of Detection and Diagnosis Methodologies in PHM Data Competitions
3.5 2023 PHM AP (Spacecraft Propulsion System)
The Japan Aerospace Exploration Agency (JAXA) initiated a competition centered on the advancement of PHM technology for spacecraft propulsion systems. The primary objective was to accurately diagnose various states ranging from normal conditions to bubble anomalies, solenoid valve faults, and unknown faults. Simulation data was sourced from four distinct spacecraft, resulting in a dataset comprising 177 training samples and 46 test samples [73]. The paucity of data posed a significant challenge for the data-driven methods.
In the competition, the champions Takanobu Minami and Jay Lee [49], introduced a novel two-step approach. Initially, a similarity measures-based model was proposed for the categorization of data into four distinct states. Subsequently, for data corresponding to the solenoid valve fault, a model incorporating physics-inspired features was employed to pinpoint the fault location and estimate the valve opening ratio. A noteworthy aspect of their methodology entailed the employment of Dynamic Time Warping (DTW) [7] on the training dataset. This step was instrumental in quantifying the variability across various segments of the sensor data. Based on data analysis, authors opted to leverage solely the initial valve opening data segment for feature extraction, thereby mitigating variations arising from individual or sample variations. This strategic refinement substantially improved the data quality, resulting in better model performance in fault detection and classification, with the team achieving 100% classification accuracy in the final competition. Standing in the second position, Lee et al. [40] devised a hybrid approach combining the XGBoost-based method and the rule-based method. While the XGBoost-based approach primarily addressed comprehensive fault classification, the rule-based method was employed to formulate the solenoid valve opening ratio equation. This was accomplished through polynomial fitting, rooted in intrinsic physical characteristics, enabling a precise estimation of the solenoid valve opening ratio. Hot on their heels, Kato et al. [29] utilized the K-NN algorithms to classify faults and pinpoint the location of anomalies. Prior to the classification, a differentiation between normal and anomalous data was executed based on a similarity method. Their estimation metrics for the valve opening ratio were hinged upon the similarity of time series waveforms. Moreover, Aimiyekagbon et al. [1] employed an ensemble framework, integrating K-means clustering and decision trees. This model, enriched by domain-specific expertise, exhibited commendable precision in both anomaly detection and fault diagnosis. A synthesis of the competition’s myriad approaches underscores the pivotal role of similarity-based methods and the extraction of physical features. These elements stand out as potent tools in the realm of fault detection and classification, especially when contending with limited datasets.
3.6 2023 IEEE (Gearbox)
Multiple fault classification under different operating conditions using vibration signals. Gearbox malfunctions in industrial machinery can result in significant downtime and damage, leading to substantial financial losses. Early detection and accurate diagnosis of these faults are paramount for ensuring the machinery’s safe operation, averting catastrophic failures, and minimizing maintenance costs. The objective of this competition [5] is to develop ML-based models that can efficiently detect faults in the planetary gearboxes of industrial machinery using vibration signals. The dataset [16] spotlights four prevalent sun gear faults: surface wear, chipped teeth, cracks, and missing teeth. Vibration signals have been recorded for a duration of five minutes each, with a sampling rate of 10 kHz, under two distinct operational conditions. Given the expansive data size, totaling 50,000 entries, the potential for deploying deep learning techniques is evident.
Various data-driven methodologies have been employed in this competition. For instance, Lee et al. [41] utilized ensemble-based CNN techniques to address fault detection via time-series vibration data. Their findings underscored the robust performance of integrating three convolution kernel-based methods such as ROCKET (RandOm Convolutional KErnel Transform) method [13], one-dimensional convolutional neural networks integrated with ResNet, and a fusion of LSTM and FCN in delivering classification results for multivariate time-series data. Similarly, both Shen et al. [60] and Matthias Kreuzer alongside Walter Kellermann [33], designed models based on residual-based CNNs techniques to navigate the fault classification challenge. Shen [60] incorporated the Short-Time Fourier Transform (STFT) [23], transforming frequency domain signals into time-frequency domain signals using Fourier analysis - an invaluable tool for interpreting time-evolving signals. On the other hand, Kreuzer and Kellermann [33] leveraged data augmentation and regularization techniques, enabling model construction with fewer parameters without sacrificing performance quality. In another noteworthy contribution, Sobha et al. [64] harnessed a tree classifier to select valuable features from raw vibration signals, subsequently developing a sequential neural network model tailored for the concurrent detection of multiple gear faults.
4 Assessment & Prognosis
This section delves into the health assessment and prognosis problems within PHM, emphasizing their role in evaluating the current health state of industrial machinery and estimating their remaining useful life. Assessment and prognosis are pivotal to implementing predictive maintenance strategies and preemptively addressing system degradation, thereby reducing the incidence of unplanned downtime, extending equipment lifespan, and optimizing maintenance schedules. Similar to the structure of Section 3, we first summarize the challenges in the competition as outlined in Table 5. Subsequently, we detail the specifics of each competition, including the data-driven approaches applied, and provide a consolidated summary of both the problems with their associated datasets and the proposed solutions in Table 6 and Table 7, respectively.
4.1 Challenges
The pivotal challenge across all competitions in health assessment and prognosis is summarized in Table 5. The common challenge focuses on identifying degradation trends and estimating RUL. In the meantime, diverse industrial scenarios present each competition with unique challenges; for instance, the 2018 PHM NA and 2020 PHM EU are both confronted with variability due to differing operational conditions—sensor discrepancies due to varied tool use in the former, and twelve distinct operational conditions influenced by solid ratios and particle sizes in the latter. Similarly, in 2021 PHM NA, the diverse flight envelopes introduce considerable variability into the data. A common thread in the 2018 and 2021 PHM NA competitions is the critical need for models to robustly classify multiple fault modes to enable precise RUL prediction. Complicating this need is the data imbalance evident in the 2018 PHM NA, where rare fault instances are overshadowed by prevalent healthy data, potentially reducing the models’ capabilities for rare fault detection. In the 2019 PHM NA, the limited availability of ultrasonic signal data, particularly for the validation set, poses a risk of model overfitting, potentially reducing the reliability of data-driven prognostics. The 2020 PHM EU is faced with the uncertainty of unknown faults, as the test set includes operational conditions not represented in the training and validation data, a discrepancy that could lead to diminished model performance. These multifaceted challenges necessitate the development of sophisticated, adaptable machine learning models capable of overcoming data discrepancy, data imbalance, and operational condition variability to deliver accurate and reliable health assessments and prognostic predictions.
Unique Challenges
2018 PHM NA
Data Discrepancy & Data Imbalance
Different Operating Conditions
Multiple Fault Modes
2019 PHM NA
Variable Loading Conditions
Limited Physical Information
Limited Data
2020 PHM EU
Different Operation Conditions
Unknown Fault
2021 PHM NA
Multiple Fault Modes
High Variability in Flight Envelopes
Common Issues: Degradation Trend Detection
Table 5: Overview of Assessment and Prognosis Challenges in PHM Data Competitions
2018 PHM NA 2019 PHM NA 2020 PHM EU 2021 PHM NA
System Ion Mill Etching System Fatigue Crack Filtration System Turbofan Engine
Failure Mode 3 1 1 7
Asset 20 8 NA NA
Operating Condition Multiple Variable Loading Conditions 12 4
Data Type Time Series Time Series Time Series Time Series
Volume of Data Limited Samples Limited Samples Limited Samples Limited Samples
(20 Train, 5 Test) (74 Train, 36 Test) (24 Train, 8 Validation, 16 Test) (90 Train Units, 38 Test Units)
Sampling Rate 0.25 Hz 5 Hz 10 Hz 1 Hz
Sampling Interval Above 70 million seconds 14000-100774 cycles 200-350 s 1-3 hours, 3-5 hours, Above 5 hours
Table 6: Overview of Assessment and Prognosis Problems in PHM Data Competitions
4.2 2018 PHM NA (Ion Mill Etching System)
In the quest to advance predictive maintenance strategies for ion mill etching systems within semiconductor manufacturing, the diagnosis and prognosis of three principal failure modes—Flowcool leak (F1), Flowcool Pressure Too High Check Flowcool Pump (F2), and Flowcool Pressure Dropped Below Limit (F3)—are critical. Successful prediction of these failures helps in the timely scheduling of maintenance operations, thereby reducing downtimes and enhancing operational efficiency. The 2018 PHM North America data challenge emphasized the analysis of fault behavior within the ion mill etch tool, tasking participants to develop a model from the sensor-derived time series data capable of accurately detecting, diagnosing, and prognosticating the time-to-failure for these malfunctions.
Investigation into the proposed solutions reveals a pronounced inclination towards tree-based methodologies for fault detection and diagnosis. Wu et al. [77], Huang et al. [22], Zhao et al. [79], and the team led by He [19] have each proposed strategies based on random forest algorithms for early degradation mode detection and diagnosis. Extending the exploration of machine learning techniques, Singh et al. [61] evaluated an assortment of models, including Logistic Regression, Generalized Linear Models, Multivariate Adaptive Regression Splines, Support Vector Regression, Random Forest, Multi-Layer Perceptron, and Gradient Boosting. Among these, the Random Forest model distinguished itself with superior performance. Zheng et al. [81], meanwhile, utilized a knowledge distillation approach towards fault detection across various modes, aiming to bolster the detection of infrequent yet critical faults.
Recent advancements in RUL prediction [30] have predominantly hinged on the application of deep-learning algorithms to analyze complex multivariate time series data. Wu et al. [77], Huang et al. [22], and Anqi He and Xiaoning Jin [19] have all utilized Long Short-Term Memory (LSTM) networks and their variations, such as Gated Recurrent Units (GRU) and LSTM-based Metric Regression (LSTM-MR), to capture important features from the raw data. Hsu et al. [20] expanded upon this approach by integrating a Temporal Convolutional Network (TCN) with LSTM, enhanced by an attention mechanism, which facilitated refined feature extraction from sensor data for accurate RUL prediction. Liu et al.’s two-stage deep transfer learning framework aimed at achieving accurate RUL prediction. In the first stage, the developed model leveraged TCN for initial temporal feature learning, followed by domain adversarial learning for data alignment based on one fault mode. Then in the second stage, the first-stage model was fine-tuned based on other fault mode data to handle multiple fault modes and enhance the RUL prediction performance [44]. Distinct from these methods, Zhao et al. [79] explored transformer networks, focusing solely on data from abnormal operation phases, while Lorenti et al. [46] provided a comprehensive comparison of state-of-the-art methods—including TCN, LSTM, attention-based mechanisms, CNN, and Transformer—culminating in the development of the CeRULEo library to facilitate research in the field. Moreover, Singh et al. [61] introduced a novel approach using Dynamic Time Warping (DTW) for RUL estimation, leveraging a library of truncated degradation curves and health score models to refine the final RUL predictions. Such diversity in methodologies not only underscores the complexity of fault diagnosis and prognosis in high-tech manufacturing but also showcases the robustness of machine learning in tackling these intricate challenges.
4.3 2019 PHM NA (Fatigue Crack)
The 2019 PHM Conference Data Challenge put a spotlight on the intricate task of fatigue crack length estimation and prediction within aluminum structures at different points. The challenge harnessed wave signal data collected by piezoelectric sensors subjected to both static and dynamic tensile stresses. Given that the test dataset was limited to wave signals from several initial loading cycles, participants were challenged to estimate crack lengths where signal data were present and predict future crack growth for specified cycles lacking signal data. The complexity of the problem delineated a clear path: data-driven methodologies were viable when signals existed, whereas scenarios lacking wave data required exploration into physics-based strategies. Reflecting this path, solutions derived from this challenge have predominantly employed a hybrid approach of data-driven and physics-based techniques.
Regarding the estimation of fatigue crack length with accessible wave signals, a variety of data-driven models and feature engineering strategies emerged, both within and beyond the competition’s scope. Karimian and colleagues [28] proposed a neural network architecture reliant on features manually derived from raw signals—such as the Pearson correlation coefficient, phase shifts, energy, and information entropy—to train models for accurate crack length estimation. On a similar trajectory, Kong and associates [32] initiated their approach with signal preprocessing, utilizing techniques like band-pass filtering and phase alignment to mitigate noise and uncertainty before applying physically insightful feature extraction methods. Thereafter, they harnessed a random forest algorithm, optimizing it through feature selection and grid search for hyperparameter fine-tuning, to estimate crack lengths. Consistent with these methods, Youn et al. [78] also leveraged a band-pass filter for feature extraction from raw wave signals, subsequently constructing a Support Vector Regression (SVR) model with hyperparameters optimized via grid search. Rao and collaborators [56], meanwhile, designed an ensemble learning regression model to improve estimation performance with four useful extracted features (root mean square value, correlation coefficient, first peak value, and logarithm of kurtosis).
Predictive modeling for scenarios lacking wave signal data required a pivot towards physics-based techniques, often in conjunction with data-driven insights, to forecast crack progression. Karimian et al. [28] devised a Particle Filter (PF) strategy, integrating the Paris Law and outputs from the previously developed neural network as observational inputs to refine and update the crack propagation pathway. Kong’s team [32] suggested an ensemble prognostics framework under consistent loading conditions, constructing a probability density function (PDF) for each instance. This was followed by a computation of weights derived from each PDF to output the final crack length prediction. Furthermore, when different loading conditions prevailed, Walker’s equation was invoked to forecast crack lengths. Innovatively, Youn et al. [78] introduced a trans-fitting approach, aimed at extracting the crack growth trend from training data and extrapolating it to the test data predictions. Additionally, Rao’s group advanced a variation version of Paris’ Law, aiming to elucidate the correlation between crack progression and the number of loading cycles [56].
4.4 2020 PHM EU (Filtration System)
In the domain of industrial maintenance, filtration systems are pivotal for helping process pollutants from industrial equipment, ensuring seamless system operation. A predominant challenge encountered within these systems is filter clogging - a phenomenon where accumulated pollutants impede flow rates, thereby disrupting standard industrial workflows. Addressing this issue, the PHM-Europe Conference 2020 data challenge concentrated on the prediction of filtration systems’ remaining useful life, crucial for timely maintenance and replacement scheduling. RUL, in this context, is delineated as the time until the pressure differential across the filter breaches a threshold of 20 psi. The challenge provided the PHME20 dataset, collected from a controlled experimental setup designed to simulate filter clogging at various contamination levels. Twelve distinct conditions were established influenced by two operational parameters: solid ratio (%) and particle size (µm). The competition predominantly witnessed the application of traditional machine-learning techniques. The champion solution [45] employed a novel hybrid approach, amalgamating kernel regression with fundamental statistical methodologies. This approach, underscored by complex data analysis, proved efficacious in enhancing the RUL prediction. Meanwhile, the runner-up, Beirami et al. [6], adopted different cutting-edge feature engineering, and machine-learning methods. Their process began with the extraction of features through a rolling window technique and proceeded with feature selection informed by the linear kernel Support Vector Machine (SVM) coefficients, Recursive Feature Elimination (RFE), correlation matrix, and monotonicity testing. A four-layer sequential neural network was their model of choice, supported by K-fold cross-validation throughout the training and validation stages. Capturing the third position, Ince et al. [26] conducted a comparative study of tree-based algorithms and the Bayesian approach. Specifically, random forest and gradient boosting and Gaussian process (GP) regression were utilized to estimate the RUL of the filtration system, supplemented by a novel fault-based RUL assignment which integrated Piecewise RUL_Assignment and Linear RUL_Assignment,
Contrasting with the feature engineering and conventional machine learning strategies, some researchers explored some deep learning methods. Vu et al. [75] introduced a Convolutional Neural Network (CNN)-based deep learning methodology for RUL prediction, marked by two architectural innovations: a Parameterized Fully Connected Layer that adjusts network weights in response to operational parameter shifts, and a multi-head predictor tailored to distinct degradation process stages. Simultaneously, Tian et al. [72] presented a new transfer ensemble learning (TEL) framework, leveraging metric learning alongside domain dissimilarity metric and Kullback–Leibler (KL) divergence, to enhance model generalization from source to target domains. This TEL framework amalgamated with a bidirectional long short-term memory (Bi-LSTM) algorithm, coined as TEL-Bi-LSTM, was offered for RUL estimation under different operating conditions. In another innovative approach, Ince and Genc [25] proposed a joint autoencoder-regression network, a deep neural architecture that fused a convolutional neural network autoencoder with a long-short-term memory network regressor in an end-to-end training paradigm. Genetic algorithms were instrumental in optimizing hyperparameters for this architecture. Additionally, Lee’s team [39] developed a distinct strategy by first establishing a health assessment criterion. They defined a Health Index (HI) for the filter system and utilized K-means clustering for the categorization of the system’s health stages. Subsequent HI predictions were facilitated by the Bi-LSTM algorithm, thus determining the system’s RUL. Lastly, Falcon et al. introduced an innovative sequence modeling technique termed the Neural Turing Machine (NTM) [17]. Conceptualized as a computational architecture, the NTM leverages available data to interact with an external memory component, an approach that facilitates enhanced accuracy in predictions. This model stands out for its ability to generate more precise outcomes when benchmarked against the prevalent LSTM-based solutions that dominate the field.
4.5 2021 PHM NA (Turbofan Engine)
The 2021 North America PHM Data Challenge was primarily focused on the prediction of RUL for turbofan engines [8], specifically under four distinct flight conditions and seven unique failure modes. Participants were required to create predictive models leveraging the N-CMAPSS dataset, aiming to accurately forecast RUL using complex condition monitoring data. The dataset, a collection consisting of 90 synthetic run-to-failure trajectories for training and an additional 38 truncated datasets for testing, served as a comprehensive foundation for developing robust predictive algorithms to predict RUL accurately.
During the competition, the winning team, led by Lovberg, put forward an innovative approach leveraging a deep convolutional neural network (DCNN) [47]. This network was distinguished by its use of dilated convolutions complemented by gated linear unit activations and integration of residual skip connections. Those techniques were designed to expand the network’s receptive field and enhance flexibility, so as to reduce the complexity of the neural network architecture by using less number of parameters, but still having comparable performance. Additionally, they adopted a strategic sequence sampling method, minimizing less informative samples while retaining enough degradation signals for the network’s input. This methodology distinguished their solution and got 1st place in the PHM data challenge competition. Moreover, Solis-Martin et al. and DeVol’s team pursued advancements in DCNNs as well. The former developed a two-level DCNN system where the first-level DCNN focused on extracting vital features from raw data and the second-level DCNN leverages the output from the first level to accurately estimate the remaining useful life [65]. The latter drew upon established deep learning architectures, deploying the basic blocks or modules from the VGG, GoogLeNet, and ResNet designs into their DCNN framework. This enabled a comparative analysis of model performances using a variety of well-known architectural features [14, 15].
Outside of the competition, Cohen et al. [10] addressed concerns regarding the uncertain and poor interpretability of deep learning models by integrating Principal Component Analysis (PCA) to refine time-domain feature sets and subsequently applying four supervised learning techniques, including artificial neural network, random forest, extreme random forest and XGBoost, to estimate RUL. Their innovative use of a custom loss function in conjunction with traditional ANN models got the best results in both Area Under the Receiver Operating Characteristic (AUROC) and Area Under the Precision-Recall (AUPR) metrics. Lastly, the domain of prognostics and health management has witnessed an upsurge in the application of XAI techniques to bolster the interpretability and trustworthiness of machine learning models. Various methods, including Local Interpretable LIME, SHAP, Layer-wise Relevance Propagation (LRP), Image-Specific Class Saliency Maps, and Gradient-weighted Class Activation Mapping (Grad-CAM), have been reported in recent literature [9, 11, 66]. These techniques are increasingly critical in elucidating the decision-making processes of complex models, thereby fostering greater confidence in their predictive capabilities.
Methodology
Deep Learning Conventional Machine Learning Feature Engineering
2018 PHM NA
Transfer Learning
LSTM, LSTM-MR, GRU
Transformer
CNN
TCN, TCN-LSTM, TCN-DANN
Random Forest, Gradient Boosting
Logistic Regression
Generalized Linear Model
SVR, MLP
MASR
DTW
SVR-RFE
Degradation Curve Library
2019 PHM NA NA
Random Forest, Gaussian Process
SVR
Neural Network
Ensemble Learning
Linear Regression
Grid Search
Genetic Algorithm
Paris’s Law
Walker’s Equation
Particle Filter
2020 PHM EU
Neural Turing Machine
Transfer Ensemble Learning
LSTM, Bi-LSTM, TEL-Bi-LSTM
Autoencoder-Regression Network
Deep CNN
Random Forest, Gradient Boosting
Gaussian Process
Kernel Regression
SVR
Ensemble Learning
Simple Statistics Method
Linear SVM Coefficient
Correlation Metric
RFE, Health Index
Monotonicity Test
2021 PHM NA
Deep CNN, FCN, VGG
ResNet(Residual Block)
GoogLeNet(Inception Module)
Random Forest, XGBoost
Extreme Random Forest
ANN
PCA
XAI,SHAP, LIME
Table 7: Overview of Assessment and Prognosis Methodologies in PHM Data Competitions
5 Common Challenges in PHM Data Challenge Competitions
In this section, we will discuss common challenges faced in the PHM data challenge competitions regarding data-related issues (missing data, data imbalance, and domain shift) and model-related issues (model selection, machine learning model interpretability, model robustness and generalization), and analyze relevant solutions applied to solve these challenges.
5.1 Data-related Issues
5.1.1 Missing Data
Handling missing data in machine learning, especially in the context of the PHM is significant, as incomplete data can significantly impact the model performance and prediction accuracy. The strategies to address missing data in PHM have evolved, encompassing a range of techniques from basic deletion to advanced imputation methods. This was evident in the 2021/2022 PHM EU competitions, where missing data posed significant challenges. In the fuse test bench dataset (2021 PHM EU), about 10% of the data was missing, and it was not evenly distributed across variables. Meanwhile, the SPI dataset(2022 PHM EU) had missing information in specific fields. To tackle these issues, various approaches were employed.
A straightforward strategy is the deletion of data points with missing values, such as listwise or pairwise deletion. For example, in the 2022 SPI dataset, instances having missing values in crucial identifiers like ”Panel_ID”, ”Figure_ID”, and ”Component_ID” were eliminated [68]. However, this approach can lead to the loss of valuable information to some degree. On the imputation front, several innovative methods were used in the 2021 PHM EU competition. One team applied the Last Observation Carried Forward (LOCF) method, coupled with backward filling, to address the gaps [12]. Interpolation was another technique used to estimate missing values, ensuring that cases with absent data did not skew the results. Additionally, Ramezani’s team introduced a novel approach, PLS-MV, a Partial Least Squares-based method for imputing missing values [55]. In addition to various imputation methods highlighted in the competitions, there remains scope for exploration in future research such as mean/median/multiple, KNN, regression imputation, maximum likelihood estimation, and machine learning-based approaches. These methods offer a promising avenue for further enhancing data quality and model reliability in PHM systems.
5.1.2 Data Imbalance
Data imbalance arises when the distribution of classes in a dataset is uneven. This issue is especially pronounced in PHM due to the rarity of failure events in comparison to data representing normal operations. Data imbalance can significantly undermine the performance of machine learning models, particularly in classification tasks. When trained on imbalanced data, models may become biased towards the majority class (normal operation) and may not effectively recognize the minority class (failure). This leads to poor performance in predicting failures, which may bring ineffective maintenance planning and unexpected downtimes. The fuse test bench dataset from the 2021 PHM EU competition and the printed circuit board dataset from the 2022 PHM EU competition exemplify this issue. The former displayed a stark imbalance with only 3-4 experiments representing each fault class against 70 experiments for the healthy class. Similarly, in the printed circuit board dataset, a staggering 95% of the SPI data was classified as healthy, highlighting a significant imbalance. The 2018 PHM NA competition also faced this challenge in the Ion Mill Etching (IME) dataset, where the faulty data is much less than normal operation data.
To counteract data imbalance in PHM, various strategies have been adopted. Resampling techniques are commonly used in competitions to adjust the dataset to balance the class distribution, either by oversampling the minority class or undersampling the majority class. Synthetic data generation is another approach, as demonstrated by a team in the 2021 PHM EU competition that used the Synthetic Minority Over-sampling Technique (SMOTE) to create synthetic samples of the minority class to balance the dataset [12]. Advanced algorithms also play a crucial role in addressing data imbalance. For instance, Tang’s team in the 2022 PHM EU competition used the Feeding Imbalance Ratio (FIR) to control the imbalance ratio in each mini-batch during neural network training [70]. Liu et al. employed transfer learning and domain adaptation techniques in the 2018 PHM NA competition to facilitate knowledge transfer across different fault modes, addressing the issue of insufficient data in specific faults [44]. Moreover, during the 2018 PHM NA competition, many teams utilized Random Forest, an ensemble learning method known for its proficiency in handling imbalanced data, by constructing a forest of decision trees. These approaches collectively demonstrate the diverse and innovative methods employed to mitigate the challenges posed by data imbalance in the PHM field.
5.1.3 Domain Shift
Domain shift refers to the changes in the data distribution between the training phase (source domain) and the real-world application phase (target domain) of the machine learning models. This phenomenon frequently occurs when models, initially trained on data from a specific set of machines or under certain conditions, are subsequently applied to different machines or varied operating conditions. Such a shift can markedly affect the performance and reliability of machine learning models in PHM, making tackling with domain shift problem essential for maintaining the robustness and effectiveness of PHM systems. Transfer learning has emerged as a primary solution to domain shift challenges, as it can achieve the adaptation of models from one domain to be effective in another through fine-tuning and domain adaptation techniques. Additionally, robust modeling approaches and the ability to continuously update models with new data are also useful and imperative for mitigating the impact of domain shift.
A clear example of domain shift was observed in the rock drill dataset from the 2022 PHM NA competition. The dataset encompassed training data from five rock drill machines, validation data from another machine, and test data from two additional machines, highlighting potential domain shift issues between training, validation, and test data. Similarly, the 2020 PHM EU dataset demonstrated domain shift, with training data featuring small and large particle sizes, while the test data included medium particle sizes. To address these challenges, various strategies were employed. In the 2022 PHM NA competition, Kim et al. utilized a domain adversarial neural network along with the minimization of maximum mean discrepancy (MMD) to tackle domain discrepancies [31]. Meanwhile, Oh et al. employed the deep CORAL method, calculating coral loss to diminish domain discrepancy effects by aligning the covariance of the six training domains with that of the test domain [53]. For the 2020 PHM EU competition, Tian’s team developed a novel Transfer Ensemble Learning (TEL) framework, facilitating knowledge transfer from the source to the target domain [72]. These approaches illustrate the innovative methodologies being implemented to overcome domain shift problems in PHM, ensuring that models remain accurate and reliable despite changes in their application environments.
5.2 Model-related Issues
5.2.1 Model Selection: Conventional Machine Learning or Deep Learning?
In PHM data challenge competitions, participants face the critical decision between employing conventional machine learning algorithms or opting for advanced deep learning models. This choice is influenced by various factors: the volume and quality of available data, and the complexity or nature of the research task, whether it be classification or prediction. An analysis of research papers from the past six years reveals insightful trends and preferences in model selection:
The availability of training data plays a significant role in determining the choice of modeling approach. Research teams often choose deep learning for building data-driven models when there is an abundance of training data. Conversely, in scenarios with limited training data, conventional machine learning methods, augmented by feature engineering, are more commonly employed. From Table 3 and Table 5, this trend is evident in the classification challenges of the 2022 PHM NA and 2023 IEEE competitions, where the large datasets provide opportunities to deploy deep learning approaches. However, in the 2022 PHM EU competition, despite sufficient data, the predominance of healthy class data over faulty data led teams to prefer traditional machine learning. In competitions like the 2021 PHM EU and 2023 PHM AP, where training data was around 100-200 samples, traditional machine learning methods were more viable due to deep learning’s requirement for a large volume of data. For RUL prediction tasks, despite smaller datasets (less than 100 samples), the nature of the data being a long time series allows for treating each time point or a series of time points (data window) as a separate training point, effectively enlarging the dataset. Thus, from table 7, we can find that both deep learning and conventional machine learning methods are utilized, except in the 2019 PHM NA competition, where the absence of wave data in part of the test data necessitated a combination of data-driven (conventional machine learning) and physics-based approaches.
In both classification and RUL prediction tasks, Table 4 and Table 7 show that tree-based methods (such as Decision Trees, Random Forest, and XGBoost) and ensemble learning are favored in traditional machine learning. These methods are known for their robustness, ability to handle various data types, and effectiveness in capturing non-linear relationships. Deep learning solutions often favor LSTM-based and CNN-based approaches, with LSTM’s proficiency in handling sequential, time-series data and CNN’s strength in feature extraction from multidimensional datasets. When domain shift is present in the competition dataset, transfer learning and domain adaptation methods are often proposed to adjust the model to new data distributions, leveraging pre-trained models to address limitations in training data. This adaptability ensures the model’s applicability and effectiveness under varied operational conditions, a frequent occurrence in PHM scenarios.
5.2.2 Machine Learning Model Interpretability
In PHM data challenge competitions, accuracy in prediction and classification is still the primary goal. However, with the advance of deep learning, emphasizing the importance of model interpretability is becoming as crucial as their accuracy in prediction and classification due to its significant impact on safety, operational decision-making, and compliance with industry standards. Model interpretability not only fosters trust and facilitates collaboration between data scientists and domain experts, but it also plays a pivotal role in error analysis and model refinement, which are especially critical in the high-stakes environment of PHM.
Despite the complexity of deep learning models, often referred to as ”black boxes,” various techniques have been developed and applied to enhance their interpretability [66, 67]. Layer-wise Relevance Propagation (LRP), Gradient-weighted Class Activation Mapping in CNNs, and attention mechanisms in sequential models are some of the methods utilized to shed light on model decision-making processes. Additionally, model-agnostic methods like LIME and SHAP have been instrumental in offering insights into model behavior. From a broader prospect, the challenge in PHM is to balance the trade-off between model accuracy and interpretability. Achieving this balance is not just about enhancing model performance, but also about building trust and ensuring compliance with industry standards. Furthermore, in an era increasingly focused on ethical and responsible AI, transparent and interpretable models are key to not only enhancing the technical aspects of AI solutions in PHM but also extending to ensuring their successful integration and acceptance in real-world applications.
5.2.3 Model Robustness and Generalization
In the realm of PHM, model robustness and model generalization are critical factors that determine the effectiveness of a model in real-world applications. Model robustness is defined as the ability of a model to maintain its performance when facing diverse challenges like noise, outliers, and adversarial examples. Techniques such as data augmentation, where training data can be expanded by creating modified versions of existing data or synthesizing new data, have become commonplace. For instance, in the 2023 IEEE competition, Kreuzer’s team employed methods like additive white Gaussian noise, circular shift, and random amplitude scaling to increase the volume of training data [33]. Moreover, regularization techniques like L1 and L2 regularization have been instrumental in preventing overfitting and enhancing stability, making the model less sensitive to small fluctuations in input data. What’s more, Ensemble methods have been increasingly recognized for their significant contribution to model robustness in PHM. A comprehensive analysis of recent studies shows that out of sixty research papers, nine utilized various ensemble techniques - including ensemble LSTM, decision trees, random forest, XGBoost, soft voting ensemble, transfer ensemble learning, ensemble regression, and ensemble CNN-based approaches - across six different PHM data challenge competitions. This trend underscores the effectiveness of ensemble methods in bolstering model robustness, particularly by reducing the influence of noise and outliers, thereby enhancing the reliability of predictive models in PHM applications. Additionally, adversarial training, which involves training models on both regular and adversarial data, has been recognized for its potential to fortify models against adversarial attacks.
Alongside robustness, model generalization has also been a focal point in PHM, with the goal being to develop models that perform reliably on new, unseen data. One key technique for improving generalization is cross-validation, which involves dividing the training data into several subsets and each time using one of the subsets as the validation data and others as the training data for model validation during the training process. In a notable example from the 2021 PHM NA competition, researchers opted for k-fold repeated random subsampling validation. This approach was chosen as an alternative to standard k-fold cross-validation to address its limitation, wherein the size of the validation set diminishes as the number of folds (k) increases [65]. In addition to cross-validation, transfer learning and domain adaptation methods have been crucial in maintaining model effectiveness and generalization across varied data distributions. These methods are particularly relevant in scenarios characterized by ”domain shift,” where the data from the target domain differs from the source domain. The specific implementations of these methodologies in various competitions have been detailed in the earlier subsubsection titled “Domain Shift”. For deep learning models, specific techniques have been increasingly adopted to improve model generalization and avoid overfitting. These include 1) dropout, where randomly selected neurons are ignored during the training stage; 2) batch normalization, which stabilizes and accelerates training by normalizing layer inputs; 3) hyperparameter optimization, where parameters such as learning rate, number of layers, and batch size are fine-tuned for optimal performance; 4) early stopping, a method where training is halted as soon as the model’s performance on a validation set begins to decline. In summary, these strategies play a critical role in enhancing the robustness and generalizability of machine learning models in the PHM domain.
6 Discussion and Prospects
Our comprehensive analysis of machine learning and deep learning methods across nine open-source industrial datasets has revealed the transformative potential that ML approaches can have on PHM. The nature of the competition tends to prioritize solutions that chiefly enhance accuracy, potentially at the expense of systematic approach, reusability, and methodological inheritance. However, the exploration of methodologies extends outside of the competition. This includes extensive research and analysis of open-source datasets, leading to the development of advanced ML and DL techniques, not just for competitive purposes but for benchmarking and comparative analysis. These efforts have notably augmented the systematization, interpretability, and applicability of ML and DL approaches, thereby expanding their utility beyond the confines of the competition-centric paradigm. Such applications have the power to bridge the gap between academic research and industrial practice, expediting the integration of data-driven machine learning for broader applications across additional industrial datasets.
Nevertheless, there are still several pivotal suggestions that merit deeper investigation and exploration by the research community going forward.
(1) Advancements in Model Interpretability, Robustness, and Generalization. The adoption of deep learning in PHM has underscored the need for models that are not only high-performing but also interpretable. Techniques such as advanced data visualization and explainable artificial intelligence methods are emerging as key tools in explaining the outputs of machine learning models. These methods are anticipated to provide industries and academia with clearer insights into the decision-making processes of PHM models, thereby building trust and facilitating more informed decision-making. Alongside interpretability, the robustness and generalization of ML models in diverse and unpredictable industrial settings are paramount. Novel approaches in transfer learning and domain adaptation can be further developed to ensure models are resilient to data variability and operational uncertainties and capable of adapting to new, unseen scenarios. This ensures their applicability in a wide range of PHM applications, driving the field towards more efficient, reliable, and transparent systems.
(2) Development of Integrated ML Toolbox. The creation of an integrated ML toolbox is essential for streamlining research in PHM. By consolidating various machine learning methods into a unified toolbox, researchers can quickly access and utilize existing tools, thus accelerating the pace of research and enhancing problem-solving efficiency. Such a toolbox allows for data analysis and ML model construction in a timely manner when encountering familiar problems, leading to substantial time savings. Moreover, the toolbox is not static; it can be continuously updated with new processing ideas and solutions, fostering an environment conducive to the rapid deployment of ML in PHM. However, this requires a collaborative effort across the PHM community.
(3) Construction of Large Knowledge Library. There have seen increasing applications of machine learning in PHM, but the knowledge and solutions remain fragmented. Therefore, there is a need to establish a comprehensive large knowledge library that can include open-source industrial sensor data, problem-solving approaches, open-source code, model weights, etc. This dynamic knowledge library could evolve over time, becoming more comprehensive and thereby benefiting the broader PHM community. For instance, Lee’s team proposed to construct a industrial large knowledge library including human-interpretable data and structured machine-generated data [38]. The proposed industrial LKL could serve as a domain-specific foundation for training Large Knowledge Models to solve domain-specific problems.
(4) Potential Utilization of Large Language Models (LLMs) and Industrial Large Knowledge Models (ILKMs). Recent advancements in large language model technologies have shown remarkable abilities in natural language processing and related tasks, hinting at the potential for general artificial intelligence applications [80]. In the PHM domain, leveraging these cutting-edge technologies could lead to groundbreaking changes. Yang et al. introduced a novel benchmark dataset focused on Question Answering (QA) in the industrial domain and proposed a new model interaction paradigm, aimed at enhancing the performance of LLMs in domain-specific QA tasks [76]. This approach signifies a substantial stride in customizing LLMs for more specialized, industry-oriented applications. Meanwhile, Li’s team systematically reviewed the current progress and key components of ChatGPT-like large-scale foundation (LSF) models, and provided a comprehensive guide on adapting these models to meet the specific needs of PHM, underscoring the challenges and opportunities for future development [42]. Moreover, as shown in Figure 3, Lee’s team proposed an Industrial Large Knowledge Model (ILKM) framework that aims to solve complex challenges in intelligent manufacturing by combining LLMs and domain-specific knowledge [38]. In summary, by integrating specialized domain knowledge with LLM technology, ML models could offer enhanced performance in addressing industrial data challenges specific to PHM.
Figure 3: Industrial Large Knowledge Model Framework [38].
7 Conclusion
In summary, machine learning gradually becomes a cornerstone in PHM, reflecting the big potential for innovative advancements in future PHM development. This paper serves as a valuable resource for both academic and industry professionals in PHM domain, offering a unified ML framework in PHM and a comprehensive overview of the current state-of-the-art machine learning approaches for diagnostics and prognostics of industrial systems using industrial open-source data. Based on two primary research task categories: ”Detection & Diagnosis” and ”Assessment & Prognosis”, we provide a detailed explanation of the problems, tasks, challenges, and relevant machine learning approaches to each competition. Furthermore, we summarize common challenges, including data-related issues and model-related issues, and analyze the solutions to address these challenges. Finally, we prospect several potential directions in the application of data-driven machine learning within PHM, encompassing improving the interpretability, robustness, and generalization of ML models, developing integrated ML toolboxes, constructing a industrial large knowledge library, harnessing the potential of LLMs and ILKMs.